OpenStudy (anonymous):

assume that $P is invested at 7% c.c. becomes $100 in a year USE A=Pe^rt a) Principal b) Amount

OpenStudy (anonymous):

i have already started the problem i have so far 100=Pe^(.07) which is 100/p=e^(.07)